OpenStudy (anonymous):

Please someone fill this out it needs to be turned in today in two hours and i don't know how the file will be in the replys

OpenStudy (anonymous):

Heres the file

OpenStudy (anonymous):

Heres the file

OpenStudy (texaschic101):

1. square A = s^2 = 6^2 = 36 P = 4s = (4)(6) = 24 2. Rhombus sorry thats all I can do now...I have to go..kinda lost track of time